OpenStudy (anonymous):

Help... Elizabeth has two identical number cubes. Both cubes have faces numbered 1 thourgh 6. if Elizabeth rolls each cube once, what is the probability that the sum of the two numbers on the top faces will be 10. Choices are 1/36 1/12 1/10 1/9

OpenStudy (anonymous):

@ospreytriple

OpenStudy (anonymous):

Hope you don't mind

OpenStudy (anonymous):

Not at all. How many different combinations are there that add up to 10?

OpenStudy (anonymous):

5+5=10 9+1=10 8+2=10 7+3=10 6+4=10

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

But each cube is numbered 1 through 6, so you can't use any numbers greater than 6

OpenStudy (anonymous):

Oh yeah

OpenStudy (anonymous):

5+5=10 6+4=10

OpenStudy (anonymous):

Umm.. Is there more? I'm not sure...

OpenStudy (anonymous):

Right. But here's a tricky bit about probability and counting. You also have to consider the order that the numbers come up. So the first cube could say 4 and the second cube could say 6. BUT the first cube could say 6 and the second cube could say 4. They gave to be counted as two different outcomes.

OpenStudy (anonymous):

So how about 5+5 6+4 4+6 Understand?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok. Now how many possible outcomes are there. Remember that each cube has six possible outcomes. How many possible outcomes are there. Remember the fundamental counting principle

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

I'm not sure...

OpenStudy (anonymous):

Ahh.. Wait so 1

OpenStudy (anonymous):

Because 4+6 and 6+4 are still the same right

OpenStudy (anonymous):

Here's a hint: 1 & 1 1 & 2 1 & 3 up to 6 & 4 6 & 5 6 & 6

OpenStudy (anonymous):

I still don't get it

OpenStudy (anonymous):

OK remember earlier that, when you had two events, you multiplied the possibilities together?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Well, there are six possibilities on the first cube and six possibilities on the second cube. How many possibilities in total?

OpenStudy (anonymous):

36

OpenStudy (anonymous):

so is it 1/36?

OpenStudy (anonymous):

Correct. And three of those possibilities give a sum of 10. What's the probability?

OpenStudy (anonymous):

5+5?

OpenStudy (anonymous):

There are 3 successful outcomes out of a total of 36 outcomes. What's the probability?

OpenStudy (anonymous):

@Madi_Lee

OpenStudy (anonymous):

3/36 x/36?

OpenStudy (anonymous):

Just 3/36. Now you have to reduce it to lowest terms

OpenStudy (anonymous):

which is 1/36

OpenStudy (anonymous):

Nope. Divide both numerator and denominator by 3 and what do you get?

OpenStudy (anonymous):

\[\frac{ 3 }{ 36 } = \frac{ 3\times1 }{ 3\times12 }=?\]

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

Oh nevermind 1 minute

OpenStudy (anonymous):

1/12

OpenStudy (anonymous):

Exactly. Well done.